OpenStudy (anonymous):

How do you simplify .7^3

OpenStudy (anonymous):

7^3 = 7x7x7 49x7 = 343

OpenStudy (anonymous):

Is it the same if it has a decimal in front of the 7?

OpenStudy (anonymous):

.7 is 7/10, then\[\left( \frac{7}{10} \right)^3=\frac{7^3}{10^3}=\frac{343}{1000}\]

OpenStudy (anonymous):

jeekim just misread the question; didn't see the decimal

OpenStudy (anonymous):

Thank you... Helping daughter with homework :) i wasnt much help so thank you!

OpenStudy (anonymous):

Sorry, didnt see the decimal Dx. It would be .343

OpenStudy (anonymous):

np

OpenStudy (anonymous):

What if its a negative number example: -7^4

OpenStudy (anonymous):

-7^4 is different from (-7)^4; -7^4=-2401 and (-7)^4=2401

OpenStudy (anonymous):

If the degree is odd, it's negative. If even, it's positive. But if the negative sign is not in a parenthesis sign, it's a negative.

OpenStudy (anonymous):

Ok thank you again!

OpenStudy (anonymous):

npa